# Continuous Optimization for Feature Selection with Permutation-Invariant Embedding and Policy-Guided Search

### v2

@@ 1. Introduction
Existing works can be divided into three categories: 1) filter methods (Yang and Pedersen, 1997; Forman et al., 2003; Hall, 1999; Yu and Liu, 2003) rank features based on specific scoring criteria. They often select the top-K ranking features as the optimal subset; 2) wrapper methods (Yang and Honavar, 1998; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997) iteratively generate and evaluate feature subsets. They select the feature subset that performs best on a downstream task as the optimal one; 3) embedded methods (Tibshirani, 1996; Sugumaran et al., 2007) integrate a feature selection regularization term (e.g., $L_{1}$ regularization) into model training. As the model converges, irrelevant features are penalized, and the non-zero coefficients indicate the optimal subset. While prior work has achieved notable success, it falls short in capturing complex feature interactions and adapting feature selection knowledge to diverse and dynamic scenarios.
@@ 4.1. Experimental Setup
We compare our method with 12 widely-used feature selection methods: (1) K-Best (Yang and Pedersen, 1997) selects top-K features with the best feature scores; (2) mRMR (Peng et al., 2005) aims to select a feature subset that comprises of the features with greatest relevance and the least redundancy to the target; (3) LASSO (Tibshirani, 1996) conducts feature selection by regularizing model parameters, that is shrinking the coefficients of useless features into zero; (4) RFE (Granitto et al., 2006) selects features in a recursive fashion, that is removing the weakest features until the stipulated feature numbers is reached; (5) LASSONET (Lemhadri et al., 2021) conducts feature selection by training neural networks with novel objective function; (6) GFS (Leardi, 1996) conducts feature selection via genetic algorithms that is recursively generating a population based on a possible feature subset first, and then using a predictive model to assess it; (7) SARLFS (Liu et al., 2021a) utilizes an agent to select features among all features to alleviate computational costs and views feature redundancy and downstream task performance as incentives; (8) MARLFS (Liu et al., 2019) is an enhanced version of SARLFS, that is building a multi-agent system to select features with the same incentive settings with SARLFS, and each agent is associated with a single feature; (9) RRA (Seijo-Pardo et al., 2017) is a rank integration algorithm that first gathers several selected feature subsets, and then integrates them based on the their rank in terms of statistical information; (10) MCDM (Hashemi et al., 2022) first combines the ranks of features from different feature selection algorithms to form a decision matrix, and then assigns scores to features based on this matrix for the purpose of more comprehensively and accurately selecting the most valuable features through considering multiple evaluation criteria; (11) GAINS (Xiao et al., 2023c) first constructs a embedding vector space by jointly optimizing an encoder-evaluator-decoder model, and then uses the gradient from well-evaluator moves the embedding point toward high performance direction to search the optimal feature subset embedding point. (12) MEL (Wang et al., 2024) divides the parent population into two sub-populations that search for optimal feature subsets independently and uses the mutual influence and knowledge sharing between the sub-populations, integrating multi-task learning with evolutionary learning. Among the introduced baseline algorithms, K-Best and mRMR are categorized as filter methods. Lasso, RFE, and LassoNet are representative of embedded methods. GFS, SARLFS, and MARLFS are considered as wrapper methods. RRA, MCDM, and GAINS are classified as hybrid feature selection methods.
@@
We conducted all experiments on the Windows 11 operating system, AMD Ryzen 5 5600X CPU, and NVIDIA GeForce RTX 3070Ti GPU, with the framework of Python 3.10.15 and PyTorch 2.5.1 (Paszke et al., 2019).
@@ 5. Related Works
Feature selection screens discriminative features from high dimensional data, improving model efficiency and performance by reducing dimensionality and shortening training time. Conventional feature selection methods are categorized into three main types: 1) Filter methods evaluate feature importance using statistical metrics (e.g., correlation, mutual information and chi-square tests) (Yang and Pedersen, 1997; Hall, 1999; Forman et al., 2003; Yu and Liu, 2003; Peng et al., 2005; Biesiada and Duch, 2008; Ding et al., 2014). However, such methods generally ignore the dependency and interactions among features. 2) Wrapper methods formulate feature selection as a search problem, training a machine learning (ML) model as the evaluation criterion and selecting the subset that yields the best performance (Liu et al., 2021b; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997; Leardi, 1996). However, such methods require excessive model training, resulting in high computational costs. 3) Embedded methods integrate feature selection into the model training process as a part of the optimization objective to evaluate the importance of features (Tibshirani, 1996; Sugumaran et al., 2007; Lemhadri et al., 2021; Granitto et al., 2006; Kumagai et al., 2022; Koyama et al., 2022). However, such methods heavily rely on the underlying model structure. Recently, generative methods have attracted substantial attention due to the success of generative AI (Xiao et al., 2023c; Ying et al., 2024a; Wang et al., 2022; Xiao et al., 2023b, a). For instance, (Xiao et al., 2023c) introduces a framework that encompasses a sequential encoder, an accuracy evaluator, a sequential decoder, and a gradient ascent optimizer, embedding discrete feature selection knowledge into a continuous embedding space and conducting gradient-ascent search process for the optimal feature subset. However, this approach is limited by: 1) incapable of encoding the permutation invariance into the embedding space; 2) the reliance on the convexity assumptions about the feature subset embedding space. To address these drawbacks, we integrate a permutation-invariant embedding mechanism with a policy-guided search strategy in CAPS. The permutation-invariant embedding is designed to mitigate the sensitivity to feature ordering, constructing a more robust embedding space that preserves feature selection knowledge. The policy-guided search leverages a reinforcement learning strategy to effectively explore this space, identifying better feature subsets without relying on strong convexity assumptions.

### v3

# arXiv:2505.11601v3[cs.LG] 27 Feb 2026

@@ 1. Introduction
Existing works can be divided into three categories: 1) filter methods (Yang and Pedersen, 1997; Forman and others, 2003; Hall, 1999; Yu and Liu, 2003) rank features based on specific scoring criteria. They often select the top-K ranking features as the optimal subset; 2) wrapper methods (Yang and Honavar, 1998; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997) iteratively generate and evaluate feature subsets. They select the feature subset that performs best on a downstream task as the optimal one; 3) embedded methods (Tibshirani, 1996; Sugumaran et al., 2007) integrate a feature selection regularization term (e.g., $L_{1}$ regularization) into model training. As the model converges, irrelevant features are penalized, and the non-zero coefficients indicate the optimal subset. While prior work has achieved notable success, it falls short in capturing complex feature interactions and adapting feature selection knowledge to diverse and dynamic scenarios.
@@ 4.1. Experimental Setup
We compare our method with 12 widely-used feature selection methods: (1) K-Best (Yang and Pedersen, 1997) selects top-K features with the best feature scores; (2) mRMR (Peng et al., 2005) aims to select a feature subset that comprises of the features with greatest relevance and the least redundancy to the target; (3) LASSO (Tibshirani, 1996) conducts feature selection by regularizing model parameters, that is shrinking the coefficients of useless features into zero; (4) RFE (Granitto et al., 2006) selects features in a recursive fashion, that is removing the weakest features until the stipulated feature numbers is reached; (5) LASSONET (Lemhadri et al., 2021) conducts feature selection by training neural networks with novel objective function; (6) GFS (Leardi, 1996) conducts feature selection via genetic algorithms that is recursively generating a population based on a possible feature subset first, and then using a predictive model to assess it; (7) SARLFS (Liu et al., 2021b) utilizes an agent to select features among all features to alleviate computational costs and views feature redundancy and downstream task performance as incentives; (8) MARLFS (Liu et al., 2019) is an enhanced version of SARLFS, that is building a multi-agent system to select features with the same incentive settings with SARLFS, and each agent is associated with a single feature; (9) RRA (Seijo-Pardo et al., 2017) is a rank integration algorithm that first gathers several selected feature subsets, and then integrates them based on the their rank in terms of statistical information; (10) MCDM (Hashemi et al., 2022) first combines the ranks of features from different feature selection algorithms to form a decision matrix, and then assigns scores to features based on this matrix for the purpose of more comprehensively and accurately selecting the most valuable features through considering multiple evaluation criteria; (11) GAINS (Xiao et al., 2023c) first constructs a embedding vector space by jointly optimizing an encoder-evaluator-decoder model, and then uses the gradient from well-evaluator moves the embedding point toward high performance direction to search the optimal feature subset embedding point. (12) MEL (Wang et al., 2024) divides the parent population into two sub-populations that search for optimal feature subsets independently and uses the mutual influence and knowledge sharing between the sub-populations, integrating multi-task learning with evolutionary learning. Among the introduced baseline algorithms, K-Best and mRMR are categorized as filter methods. Lasso, RFE, and LassoNet are representative of embedded methods. GFS, SARLFS, and MARLFS are considered as wrapper methods. RRA, MCDM, and GAINS are classified as hybrid feature selection methods.
@@
We conducted all experiments on the Windows 11 operating system, AMD Ryzen 5 5600X CPU, and NVIDIA GeForce RTX 3070Ti GPU, with the framework of Python 3.10.15 and PyTorch 2.5.1.
@@ 5. Related Works
Feature selection screens discriminative features from high dimensional data, improving model efficiency and performance by reducing dimensionality and shortening training time. Conventional feature selection methods are categorized into three main types: 1) Filter methods evaluate feature importance using statistical metrics (e.g., correlation, mutual information and chi-square tests) (Yang and Pedersen, 1997; Hall, 1999; Forman and others, 2003; Yu and Liu, 2003; Peng et al., 2005; Biesiada and Duch, 2008; Ding et al., 2014). However, such methods generally ignore the dependency and interactions among features. 2) Wrapper methods formulate feature selection as a search problem, training a machine learning (ML) model as the evaluation criterion and selecting the subset that yields the best performance (Liu et al., 2021a; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997; Leardi, 1996). However, such methods require excessive model training, resulting in high computational costs. 3) Embedded methods integrate feature selection into the model training process as a part of the optimization objective to evaluate the importance of features (Tibshirani, 1996; Sugumaran et al., 2007; Lemhadri et al., 2021; Granitto et al., 2006; Kumagai et al., 2022; Koyama et al., 2022). However, such methods heavily rely on the underlying model structure. Recently, generative methods have attracted substantial attention due to the success of generative AI (Xiao et al., 2023c; Ying et al., 2024a; Wang et al., 2022; Xiao et al., 2023b, a; Wang et al., 2026; Liu et al., 2025, 2026b, 2026a). For instance, (Xiao et al., 2023c) introduces a framework that encompasses a sequential encoder, an accuracy evaluator, a sequential decoder, and a gradient ascent optimizer, embedding discrete feature selection knowledge into a continuous embedding space and conducting gradient-ascent search process for the optimal feature subset. However, this approach is limited by: 1) incapable of encoding the permutation invariance into the embedding space; 2) the reliance on the convexity assumptions about the feature subset embedding space. To address these drawbacks, we integrate a permutation-invariant embedding mechanism with a policy-guided search strategy in CAPS. The permutation-invariant embedding is designed to mitigate the sensitivity to feature ordering, constructing a more robust embedding space that preserves feature selection knowledge. The policy-guided search leverages a reinforcement learning strategy to effectively explore this space, identifying better feature subsets without relying on strong convexity assumptions.